# Greedy Heuristics for Sampling-based Motion Planning in High-Dimensional State Spaces

### v1

@@ 6 Simulation and Experiments
Greedy RRT* was compared against the Open Motion Planning Library (OMPL; Sucan et al., 2012) implementations of RRT-Connect, RRT*, Informed RRT*, RRT*-Connect (Klemm et al., 2015), BIT* (Gammell et al., 2020) and AIT* (Strub and Gammell, 2020, 2022) on simulated planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$ (Section 6.1), manipulation problems for two Barret WAM arms with $14$ degrees-of-freedom ($\mathbb{R}^{14}$) in the Open Robotics Automation Virtual Environment (OpenRAVE; Diankov, 2010) (Section 6.2) and real-world planning problem for a self-reconfigurable mobile robot Panthera in SE$(2)$ and SE(2)+$(2)+\mathbb{R}^{2}$ state spaces (Section 6.3) [^1] Simulations were run in Ubuntu 22.04 docker on a MacBook Air with 16 GB of RAM and Apple M2 chip.. Planner developer tools (PDT; Gammell et al., 2022) was used to run all the experiments.
@@
The planners were tested with the optimization objective of minimizing path length. All the planners used maximum edge lengths of $0.3$, $0.4$, $0.5$, $0.65$, $1.25$, $2.4$ and $3.0$ in $\mathbb{R}^{2}$, SE$(2)$, $\mathbb{R}^{4}$, SE(2)+$(2)+\mathbb{R}^{2}$, $\mathbb{R}^{8}$, $\mathbb{R}^{14}$ and $\mathbb{R}^{16}$ respectively. An RGG constant of $\eta=1.001$ was used for all the planners and the Euclidean distance (i.e., $L^{2}$ norm) for path length was used as admissible cost heuristics. BIT*-based planners used a batch size of $100$ and G-RRT* used the greedy biasing ratio of $\epsilon=0.9$ to urge the exploitation in all the experiments.
6.1 Abstract planning problems
@@
The planners were tested on three simulated planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$, and $\mathbb{R}^{16}$ as depicted in Figure 6. The first problem consists of many axis-aligned homotopy classes and/or (hyper)rectangles, filled over the whole problem state space, where the initial and goal states are located at $[-0.25,0.0,\dots,0.0]^{T}$ and $[0.25,0.0,\dots,0.0]^{T}$, respectively (Figure 6(a)). The second problem contains two homotopy classes with a narrow passage gap of width $0.04$, where the initial and goal states are located at $[-0.3,0.0,\dots,0.0]^{T}$ and $[0.3,0.0,\dots,0.0]^{T}$, respectively (Figure 6(b)). The last experiment is a challenging dual-enclosure problem from planning literature, consisting of two hollow axis-aligned (hyper)rectangles enclosing the initial and goal states. Here, the initial and goal states are located at $[-0.3,0,\dots,0]^{T}$ and $[0.3,0,\dots,0]^{T}$, respectively (Figure 6(c)). Note that, the planning time given to the planners varies based on the difficulty of the problem domain and the number of state dimensions. Each planner was run $100$ trials with different pseudorandom seeds in which the solution cost was recorded every $10^{-4}$ seconds by a separate thread. If the solution is not found yet during recording, we set the value of a solution cost to be infinity for the purpose of calculating the median. Figures 7, 8 and 9 present the performance of the tested planners on these planning problems, where the median solution cost of each planner, together with its success rate is plotted versus computation time.
@@
These results demonstrate the necessity of exploitation techniques in sampling-based planners for tackling challenging planning problems. G-RRT* performs comparably to AIT* and BIT* in lower state dimensions and consistently outperforms all other RRT*-based algorithms. It finds solutions in every trial with a 100% success rate and is competitive with RRT-Connect algorithm in terms of success rate. In higher dimensions, G-RRT* outperforms all the anytime sampling-based planners, especially in converging to the optimal solutions.
@@
From the results, we can observe that the performance of the RRT-based planners decreases as the state dimension grows. These single-tree planners struggle to find solutions within the available planning time, especially in $\mathbb{R}^{16}$ for the experiments from Figures 6(a) and 6(b). Note that we omitted running single-tree RRT versions in the double enclosures experiment (Figure 6(c)) since it is a highly challenging problem, widely recognized as a bug trap problem, for any sampling-based planning algorithm in high dimensions (Yershova et al., 2005).
6.1.1 Problems with axis-aligned homotopy classes
@@
In this problem involving many axis-aligned homotopy classes (Figure 6(a)), there is not much difference in performance between the planners in $\mathbb{R}^{2}$ and $\mathbb{R}^{4}$ (Figures 7(a) and 7(b)). Here, G-RRT*, BIT*, AIT*, and RRT-Connect find solutions faster than other RRT*-based planners with a similar success rate. However, as the state dimension grows—i.e., in $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$, G-RRT* outperforms all the tested planners in converging to the lowest cost in a shorter time period, followed by BIT* and AIT* algorithms (Figures 7(c) and 7(d)). Note that in $\mathbb{R}^{16}$ problems, RRT*-based planners are not able to find solutions within the available time budget.
6.1.2 Narrow passage gap problems
@@
In this experiment involving a narrow passage gap (Figure 6(b)), all the tested planners were able to find solutions in $\mathbb{R}^{2}$ and $\mathbb{R}^{4}$ problems with a 100% success rate (Figures 8(a) and 8(b)). Here, G-RRT*, BIT*, and AIT* planners converge to the optimum at a similar rate when finding a solution through a narrow passage gap. However, RRT*-based algorithms struggle to find a path through the narrow passage gap in $\mathbb{R}^{4}$. On the other hand, there is a noticeable performance change in higher state dimensions, such as $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$ problems. All asymptotically optimal planners struggle to find a path through a narrow passage gap, but G-RRT* is the only planner that is able to find a path through it within the available time budget, and the difference becomes larger as the state dimension increases (Figures 8(c) and 8(d)).
6.1.3 Double enclosures problems
@@
We also conducted experiments on challenging problems involving double enclosures symmetric around the start and goal states. Since all asymptotically optimal sampling-based planners struggled to solve this problem as state dimensions increased, it was only run for $\mathbb{R}^{4}$ and $\mathbb{R}^{8}$ problems, and RRT*-based planners were omitted from the experiments. In $\mathbb{R}^{4}$, all the tested planners were able to solve the problem within the available time and showed comparable performance in converging to the optimum (Figure 9(a)). However, all asymptotically optimal planners struggled to solve this problem in higher state dimensions, specifically in $\mathbb{R}^{8}$. G-RRT* and RRT-Connect were the only planners that found a solution in more than 50% of the trials (Figure 9(b)). This demonstrates the importance of maintaining two greedily connecting trees in problems involving several biased Voronoi regions (Yershova et al., 2005). Nevertheless, since RRT-Connect is a non-anytime planner that cannot converge to the optimum, G-RRT* is the only almost-surely asymptotically optimal planner that requires less time to find better median solution costs within an available planning time in higher state dimensions.
6.2 Manipulation Problems
@@
The planners were also tested on manipulation tasks involving Barrett Whole-Arm Manipulator (WAM) arms within the Open Robotics Automation Virtual Environment (OpenRAVE; Diankov, 2010). OpenRAVE was set up to utilize the Flexible Collision Library (FCL; Pan et al., 2012), with a collision checking resolution set at $10^{-2}$. Figure 10 depicts the dual-arm planning experiment for two Barrett WAM arms with $14$ degrees-of-freedom. The objective is to move them from an initial configuration, where they point in the same forward direction, to a goal configuration where the arms are extended outward in opposite directions without colliding with the cage. Results for the performance of all the tested algorithms when optimizing path length are illustrated in Figure 11.
@@
From the results, it is evident that RRT* and Informed RRT* planners struggle to find solutions within the available planning time due to the high dimensionality of state spaces. In contrast, G-RRT* quickly finds initial solutions, as fast as RRT-Connect and RRT*-Connect algorithms, all achieving a 100% success rate, followed by BIT* and AIT* planners. In summary, G-RRT* outperforms all tested asymptotically optimal planners in converging to optimal solutions, while also matching the speed of non-optimal planners in finding initial solutions.
6.3 Planning for Panthera reconfigurable robot in real-world environment
@@
We validate the proposed G-RRT* algorithm with a self-reconfigurable robot, Panthera, developed by Hayat et al. (2019) as a case study. Panthera has four independent steering units designed for omnidirectional locomotion. Its body frame can also be expanded and compressed by independently controlling the additional two gaits to traverse tight spaces, as observed in Figure 12. The state of Panthera in the world frame can then be represented as a 5-dimensional space, ${x,y,\theta,j_{0},j_{1}}$, where ${x,y,\theta}$ denotes the SE$(2)$ state space, and ${j_{0},j_{1}}$ are the two controllable gaits of Panthera in $\mathbb{R}^{2}$. Here, the angles for each steering unit of Panthera are not considered during planning since this paper focuses only on geometric path planning. The resultant paths from the planners consist solely of waypoints representing the robot’s state space, and an external local trajectory controller is responsible for following these generated waypoints.
@@
Figure 12: Panthera reconfigurable pavement sweeping platform with expanded and compressed ability.
@@
Figure 14: The robot navigation trajectories (green color) in the 3D map of the real testbed workspace. (a) RRT* without reconfiguration, (b) G-RRT* without reconfiguration, (c) RRT* with reconfiguration, and (d) G-RRT* with reconfiguration.
To evaluate the efficiency of saving energy and time of different trajectories generated from a predefined source to destination inside the real-world environment, the solution paths of G-RRT* were compared against RRT* with informed sampling in both SE$(2)$ (i.e., without reconfiguration) and SE(2)+$(2)+\mathbb{R}^{2}$ state spaces (i.e., with reconfiguration) using the Panthera robot as a case study. Note that G-RRT* and RRT* were integrated into Panthera’s navigation software stack as the global path planners using the ROS framework (Quigley et al., 2009). To enable the reconfigurable ability, we set the flexible robot base width as the adaptive footprint parameter in the global planning layer. The experiments were conducted for $10$ trials, and the solutions of G-RRT* and RRT* were selected after $10$ seconds of planning time for SE$(2)$ and $30$ seconds for SE(2)+$(2)+\mathbb{R}^{2}$ to conduct the navigation experiments.
Using HDL 3D SLAM (Koide et al., 2018) with a 3D Velodyne 16-channel sensor, the 3D workspace was mapped, as shown in Figure 13. The workspace consists of several possible options (Path 1, 2, and 3, as seen in Figure 13) with different path lengths to connect the predefined source to destination waypoints. Once the global paths were generated to connect these waypoints, the robot followed the path, executing the velocity commands issued by a local trajectory controller to visit the waypoints orderly. The time spent and energy consumption were recorded by Panthera’s battery management system after each navigation. The example path of each tested method is shown in Figure 14. In the case of SE$(2)$ state space (i.e., a 3-dimensional space without reconfiguration), both RRT* and G-RRT* found the solution path through path option 1, and in the case of SE(2)+$(2)+\mathbb{R}^{2}$ state space (i.e., a 5-dimensional space with reconfiguration), both planners were able to find the path through option 2 with a narrow passage. Nevertheless, the solution paths derived by G-RRT* within the available time intervals were smoother and had fewer twists and turns than RRT*.
Table 1: Execution time and energy consumption of the planners for the real-world testbed environment.
@@
The average energy consumption and navigation time after $10$ trials are presented in Table 1. We can observe that G-RRT* outperforms RRT* in terms of both energy usage and time spent, and the margins of outperformance are significantly larger in higher-dimensional state spaces. Specifically, in the case of a 3-dimensional space without reconfiguration, the time and energy spent by RRT* are only slightly higher than those of the proposed method G-RRT*, at about 6.69% and 7.47%, respectively. However, when the search space increases to 5 dimensions, considering the reconfiguration of Panthera during planning, RRT* consumes significantly more time and energy than G-RRT*, as expected due to the increased complexity in finding optimal solutions. The difference in time and energy consumption for RRT* in this case increases to 16.35% and 11.01%, respectively, resulting in a trajectory with more turns and less smoothness.
7 Conclusion and Discussion

### v2

@@ VI Simulation and Experiments
Greedy RRT* was compared against the Open Motion Planning Library (OMPL) [43] implementations of RRT-Connect, RRT*, Informed RRT*, RRT*-Connect [21], BIT* [44] and AIT* [45, 46] on simulated planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$ (Section VI-A), manipulation problems for two Barret WAM arms with $14$ degrees-of-freedom ($\mathbb{R}^{14}$) in the Open Robotics Automation Virtual Environment (OpenRAVE) [47] (Section VI-B) and real-world planning problem for a self-reconfigurable mobile robot Panthera in SE$(2)$ and SE(2)+$(2)+\mathbb{R}^{2}$ state spaces (Section VI-D) [^1] Simulations were run in Ubuntu 22.04 docker container on a MacBook Air with 16 GB of RAM and Apple M2 chip.. Planner developer tools (PDT) [48] was used to run all the experiments.
@@
The planners were tested with the optimization objective of minimizing path length. All the planners used maximum edge lengths of $0.3$, $0.4$, $0.5$, $0.65$, $1.25$, $2.4$ and $3.0$ in $\mathbb{R}^{2}$, SE$(2)$, $\mathbb{R}^{4}$, SE(2)+$(2)+\mathbb{R}^{2}$, $\mathbb{R}^{8}$, $\mathbb{R}^{14}$ and $\mathbb{R}^{16}$, respectively. An RGG constant of $\eta=1.001$ was used for all the planners, and the Euclidean distance (i.e., $L^{2}$ norm) for path length was used as an admissible cost heuristic. The BIT*-based planners used a batch size of $100$, and G-RRT* used a greedy biasing ratio of $\epsilon=0.9$ to urge the exploitation in all the experiments.
VI-A Abstract planning problems
@@
The planners were tested on three simulated planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$, and $\mathbb{R}^{16}$ as depicted in Figure 7. The first problem consists of many axis-aligned homotopy classes and/or (hyper)rectangles, filled over the whole problem state space, where the initial and goal states are located at $[-0.25,0.0,\dots,0.0]^{T}$ and $[0.25,0.0,\dots,0.0]^{T}$, respectively (Figure 7(a)). The second problem contains two homotopy classes with a narrow passage gap of width $0.04$, where the initial and goal states are located at $[-0.3,0.0,\dots,0.0]^{T}$ and $[0.3,0.0,\dots,0.0]^{T}$, respectively (Figure 7(b)). The last experiment is a challenging dual-enclosure problem from planning literature, consisting of two hollow axis-aligned (hyper)rectangles enclosing the initial and goal states. Here, the initial and goal states are located at $[-0.3,0,\dots,0]^{T}$ and $[0.3,0,\dots,0]^{T}$, respectively (Figure 7(c)). Note that the planning time given to the planners varies based on the difficulty of the problem domain and the number of state dimensions. Each planner was run $100$ trials with different pseudorandom seeds in which the solution cost was recorded every $10^{-4}$ seconds by a separate thread. If the solution is not found yet during recording, we set the value of a solution cost to be infinity for the purpose of calculating the median. Figures 8, 9 and 10 present the performance of the tested planners on these planning problems, where the median solution cost of each planner, together with its success rate, is plotted versus computation time.
@@
These results demonstrate the necessity of exploitation in sampling-based planners for tackling challenging planning problems. G-RRT* performs comparably to AIT* and BIT* in lower state dimensions and consistently outperforms all other RRT*-based algorithms. It finds solutions in every trial, achieving a 100% success rate, and is competitive with the RRT-Connect algorithm in terms of success rate. In higher dimensions, G-RRT* outperforms all the anytime sampling-based planners, particularly in converging to optimal solutions.
@@
From the results, we can observe that the performance of the RRT-based planners decreases as the state dimension grows. These single-tree planners struggle to find solutions within the available planning time, especially in $\mathbb{R}^{16}$, as shown in the experiments from Figures 7(a) and 7(b). Note that we omitted running the single-tree RRT versions in the double enclosures experiment (Figure 7(c)) since it is a highly challenging problem, widely recognized as a bug trap problem for any sampling-based planning algorithm in high dimensions [49].
VI-A1 Problems with axis-aligned homotopy classes
@@
In this problem involving many axis-aligned homotopy classes (Figure 7(a)), there is not much difference in performance between the planners in $\mathbb{R}^{2}$ and $\mathbb{R}^{4}$ (Figures 8(a) and 8(b)). Here, G-RRT*, BIT*, AIT*, and RRT-Connect find solutions faster than other RRT*-based planners with a similar success rate. However, as the state dimension grows—i.e., in $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$, G-RRT* outperforms all the tested planners in converging to the lowest cost in a shorter time period, followed by BIT* and AIT* algorithms (Figures 8(c) and 8(d)). Note that in $\mathbb{R}^{16}$ problems, RRT*-based planners are not able to find solutions within the available time budget.
VI-A2 Narrow passage gap problems
@@
In this experiment involving a narrow passage gap (Figure 7(b)), all the tested planners were able to find solutions in $\mathbb{R}^{2}$ and $\mathbb{R}^{4}$ problems with a 100% success rate (Figures 9(a) and 9(b)). Here, G-RRT*, BIT*, and AIT* planners converge to the optimum at a similar rate when finding a solution through a narrow passage gap. However, RRT*-based algorithms struggle to find a path through the narrow passage gap in $\mathbb{R}^{4}$. On the other hand, there is a noticeable performance change in higher state dimensions, such as $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$ problems. All asymptotically optimal planners struggle to find a path through a narrow passage gap, but G-RRT* is the only planner that is able to find a path through it within the available time budget, and the difference becomes larger as the state dimension increases (Figures 9(c) and 9(d)).
VI-A3 Double enclosures problems
@@
We also conducted experiments on challenging problems involving double enclosures symmetric around the start and goal states. Since all asymptotically optimal sampling-based planners struggled to solve this problem as state dimensions increased, it was only run for $\mathbb{R}^{4}$ and $\mathbb{R}^{8}$ problems, and RRT*-based planners were omitted from the experiments. In $\mathbb{R}^{4}$, all the tested planners were able to solve the problem within the available time and showed comparable performance in converging to the optimum (Figure 10(a)). However, all asymptotically optimal planners struggled to solve this problem in higher state dimensions, specifically in $\mathbb{R}^{8}$. G-RRT* and RRT-Connect were the only planners that found a solution in more than 50% of the trials (Figure 10(b)). This demonstrates the importance of maintaining two greedily connecting trees in problems involving several biased Voronoi regions [49]. Nevertheless, since RRT-Connect is a non-anytime planner that cannot converge to the optimum, G-RRT* is the only almost-surely asymptotically optimal planner that requires less time to find better median solution costs within an available planning time in higher state dimensions.
VI-B Manipulation Problems
@@
The planners were also tested on manipulation tasks involving Barrett Whole-Arm Manipulator (WAM) arms within the Open Robotics Automation Virtual Environment (OpenRAVE) [47]. OpenRAVE was set up to utilize the Flexible Collision Library (FCL) [50], with a collision checking resolution set at $10^{-2}$. Figure 11 depicts the dual-arm planning experiment for two Barrett WAM arms with $14$ degrees-of-freedom. The objective is to move them from an initial configuration, where they point in the same forward direction, to a goal configuration where the arms are extended outward in opposite directions without colliding with the cage. Results for the performance of all the tested algorithms when optimizing path length are illustrated in Figure 12.
@@
From the results, it is evident that RRT* and Informed RRT* planners struggle to find solutions within the available planning time due to the high dimensionality of state spaces. In contrast, G-RRT* quickly finds initial solutions, as fast as RRT-Connect and RRT*-Connect algorithms, all achieving a 100% success rate, followed by BIT* and AIT*. In summary, G-RRT* outperforms all tested asymptotically optimal planners in converging to optimal solutions, while also matching the speed of non-optimal planners in finding initial solutions.
(a) Many homotopy classes in $\mathbb{R}^{2}$
(b) Many homotopy classes in $\mathbb{R}^{4}$
(c) Many homotopy classes in $\mathbb{R}^{8}$
(d) Narrow passage gap in $\mathbb{R}^{2}$
(e) Narrow passage gap in $\mathbb{R}^{4}$
(f) Narrow passage gap in $\mathbb{R}^{8}$
(g) Double Enclosure in $\mathbb{R}^{2}$
(h) Double Enclosure in $\mathbb{R}^{4}$
Figure 13: Effect of varying greedy biasing ratios on the planner’s convergence rate for the experiment described in Section VI-C. The plots show the median solution cost versus running time for the abstract planning problems outlined in Section VI-A. In these experiments, RRT* was modified to incorporate greedy informed sampling, using biasing ratio values of 00, $0.5$, $0.9$, and $1.0$. Each variation was tested over 100 trials in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, and $\mathbb{R}^{8}$ for problems involving many homotopy classes and narrow passage gap (plots a-f) and in $\mathbb{R}^{2}$ and $\mathbb{R}^{4}$ for the double enclosure problem (plots g-h).
VI-C Effect of greedy biasing ratio on convergence rate
In this section, we investigate how varying the greedy biasing ratio $\epsilon$ influences the convergence rate, exploitation vs. exploration balance, and the overall solution quality for different types of planning problems in both low and high-dimensional spaces. We experiment with values of 00, $0.5$, $0.9$, and $1.0$ to represent a spectrum from pure exploration to pure exploitation. At $\epsilon=0$, the planner performs fully informed sampling, focusing on exploration. As epsilon increases, the planner shifts toward greedy informed sampling, with $\epsilon=1.0$ representing complete exploitation.
To assess the impact of the greedy biasing ratio on convergence, we evaluate the single-tree-based RRT* algorithm with greedy informed sampling, varying the $\epsilon$ values. These variants are applied to the abstract planning problems introduced in Section VI-A, focusing on $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, and $\mathbb{R}^{8}$ spaces. We exclude $\mathbb{R}^{16}$ due to the difficulty single-tree planners face in finding initial solutions in high-dimensional spaces. For similar reasons, only $\mathbb{R}^{2}$ and $\mathbb{R}^{4}$ spaces are considered for the double enclosure problems.
Figure 13 presents the greedy exploitation effect across different planning problems and state dimensions. Planners using $\epsilon=0.9$ and $\epsilon=1.0$ outperform all other variants in converging faster to the optimum, followed by the planner using $\epsilon=0.5$ (i.e., half exploration with the informed set and half exploitation with the greedy informed set). When $\epsilon=0.0$ is used (i.e., full exploration with the informed set), it struggles to converge to the optimal solution, and this gap becomes larger as the difficulty of the planning problem increases. Notably, planners using $\epsilon=0.9$ and $\epsilon=1.0$ show similar behavior across all problems, regardless of state dimensions.
This demonstrates that an $\epsilon$ value of 0.9 is sufficient to significantly improve the convergence rate of planning algorithms using direct sampling procedures. The remaining 10% chance also preserves optimality due to the likelihood of correctly sampling the omniscient set by sampling the informed set, which is a necessary condition to improve the current solution to a holonomic problem (Lemma 12, [31]). In practice, however, this difference can be ignored, and using $\epsilon=1.0$ will still lead to convergence and preserve optimality, though there is a risk of getting stuck in local minima in certain planning problems (see Theorem 2, Section III).
VI-D Planning for Panthera reconfigurable robot in real-world environment
@@
We validate the proposed G-RRT* algorithm using a self-reconfigurable robot, Panthera, developed by [14] as a case study. Panthera has four independent steering units designed for omnidirectional locomotion. Its body frame can also be expanded and compressed by independently controlling two additional gaits to traverse tight spaces, as shown in Figure 14. The state of Panthera in the world frame can be represented as a 5-dimensional space, ${x,y,\theta,j_{0},j_{1}}$, where ${x,y,\theta}$ denotes the SE$(2)$ state space, and ${j_{0},j_{1}}$ are the two controllable gaits of Panthera in $\mathbb{R}^{2}$. Here, the angles for each steering unit of Panthera are not considered during planning, as this paper focuses only on geometric path planning. The resultant paths from the planners consist solely of waypoints representing the robot’s state space, with an external local trajectory controller responsible for following these generated waypoints.
@@
Figure 14: Panthera reconfigurable pavement-sweeping platform with the ability to expand and compress.
@@
Figure 16: The robot navigation trajectories (in green) in the 3D map of the real testbed workspace. (a) Informed RRT* without reconfiguration, (b) G-RRT* without reconfiguration, (c) Informed RRT* with reconfiguration, and (d) G-RRT* with reconfiguration.
To evaluate the efficiency of trajectories generated from a predefined source to destination inside the real-world environment, the solution paths of G-RRT* were compared against Informed RRT* in both SE$(2)$ (i.e., without reconfiguration) and SE(2)+$(2)+\mathbb{R}^{2}$ state spaces (i.e., with reconfiguration) using the Panthera robot as a case study. G-RRT* and Informed RRT* are integrated into the navigation stack of the ROS framework [51] as global planner plugins. Once the global paths were generated to connect these waypoints, the robot followed the path within the prebuilt map, executing the velocity commands issued by a local trajectory controller to track the waypoints along the global path in an orderly manner. The local path planner used in this work is based on our previous study [52], which was also integrated into the ROS framework for Panthera. To enable reconfigurability, the robot’s flexible base width was set as an adaptive footprint parameter in the global planning layer, with the configuration states adjusted at each waypoint according to the specific methods being tested. The experiments were conducted for $10$ trials, and the solutions of G-RRT* and Informed RRT* were selected after $10$ seconds of planning time for SE$(2)$ and $30$ seconds for SE(2)+$(2)+\mathbb{R}^{2}$ to conduct the navigation experiments.
Using HDL 3D SLAM [53] with a 3D Velodyne 16-channel sensor, the 3D workspace was mapped, as shown in Figure 15. The workspace consists of several possible options (Path 1, 2, and 3, as seen in Figure 15) with different path lengths to connect the predefined source to destination waypoints. The example global path of each tested method is shown in Figure 16. Among the three path options in the tested map, both Informed RRT* and G-RRT* found a solution through path option 1 in the SE$(2)$ state space (i.e., a 3-dimensional space without reconfiguration). In the SE(2)+$(2)+\mathbb{R}^{2}$ state space (i.e., a 5-dimensional space with reconfiguration), both planners found a solution through path option 2, which involved a narrow passage. However, the paths generated by G-RRT* within the available time intervals were approximately 9% shorter and exhibited fewer twists and turns compared to those produced by Informed RRT*.
TABLE I: Execution time and energy consumption of the planners for the real-world testbed environment.
@@
A smoother path allows the robot to follow the trajectory more easily, contributing to improved energy efficiency. Hence, in addition to path length, numerical comparisons between G-RRT* and Informed RRT* also considered the time spent and energy consumption during navigation. Energy consumption data were obtained from Panthera’s battery management system, with the current sampled at $10$ kHz at $48$ V. The maximum motor speed during the experiments was $150$ rpm. By integrating these current measurements over time, the total energy consumption can be calculated while the robot navigates the path. The average energy consumption and navigation time after $10$ trials are presented in Table I. We can observe that G-RRT* outperforms Informed RRT* in terms of both energy usage and time spent, with the margins of outperformance being significantly larger in higher-dimensional state spaces. Specifically, in the case of a $3$-dimensional space without reconfiguration, the time and energy spent by Informed RRT* are only slightly higher than those of the proposed method, G-RRT*, at about $6.69$% and $8.08$%, respectively. However, when the search space increases to $5$ dimensions, considering the reconfiguration of Panthera during planning, Informed RRT* consumes significantly more time and energy than G-RRT*, as expected due to the increased complexity in finding optimal solutions. The difference in time and energy consumption for Informed RRT* in this case increases to $19.56$% and $12.36$%, respectively, resulting in a trajectory with more turns and less smoothness.
VII Conclusion and Discussion